OpenStudy (anonymous):

Find the edge length of a cube if the surface area is 96cm^2?

OpenStudy (anonymous):

The total surface area is 6 times the surface area of each face, that is: \[A = 6 * l^2\] Where A is the total surface area and l is the side length of the cube. Now: \[A = 6 * l^2 = 96 cm^2\]So rearranging the formula gives: \[l = \sqrt{\frac{96}{6}}\]To give an answer of \[l = \sqrt{16} = 4\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome.